Curriculum Vitae

David Bryan Carpenter

January 10, 1992

Personal Details

Date of Birth: 15th January, 1958

Place of Birth: Birmingham, UK

Marital Status: Single

First Degree: BSc., Physics, from Westfield College, London.

First Class Honours. Degree awarded 1979.

Second Degree: PhD., Theoretical High Energy Physics, also from Westfield. Supervisor: Prof. E. Leader. Degree awarded 1984.

Thesis: Lattice and Large N Studies of Two-dimensional QCD, submitted to University of London, 1983.

Positions:

1983-1984 Research Fellow in Theoretical Physics Department, Edinburgh University.

1985 Held Royal Society Overseas Fellowship at DESY, Hamburg.

1985-1988 Research Fellow in Computational Physics Group, Southampton.

1989 Employed at “Transputer Technology Solutions”, Southampton.

1989-91 Research Fellow in Department of Electronics and Computer Science, Southampton (funded by Esprit Project PUMA).

1991-92 Research Fellow in Department of Electronics and Computer Science and Department of Physics, Southampton.

Research Interests

After spending a couple of years in my early research career involved in more or less theoretical work on lattice gauge theories (and two-dimensional field theory), I have for the last five years been very heavily involved in computational work, programming parallel computers, most especially transputers, to simulate physical systems.

My doctoral research involved applying the at that time newly-developed computational techniques for handling lattice fermions to a low-dimensional field theory [1], together with application of a novel analytic technique to the same theory. Subsequently at Edinburgh I was exposed to one of the early successful massively parallel computers— the ICL DAP, although at that stage, and for the next year at DESY, my main interests remained more analytic [2] [3] [4].

In 1985 I came to Southampton to work with Prof. A.J.G. Hey, then in the Physics Department, on the earliest transputer systems. In close collaboration with inmos we wrote some of the first scientific programs to exploit this new type of machine, at that stage concentrating on statistical physics problems. In particular we worked with J.T. Chalker on elucidating the phase diagram of a modified XY model of theoretical interest at that time [5] [8].

During the ESPRIT supernode project, which developed a reconfigurable multitransputer supercomputer, I worked on programming paradigms for exploiting transputer networks in simulation of physical systems [6]. Later on this experience was applied to QCD simulations, for example contributing simulation results obtained on a Meiko transputer machine to a study of glueballs with C. Michael and M.J. Teper [9].

At the end of the supernode project Southampton was provided with a 128 floating-point-transputer machine. In the last year this machine (which, with careful programming, is capable of CRAY-like performance) was utilised to perform a calculation of heavy meson decay constants in a collaboration with Prof. Sachrajda [10]. Currently we are attempting to perform a more accurate calculation, in a collaboration involving contributions from several European groups and several machines [11].

My experience over several years now has given me a strong grounding in exploiting parallel computers to the maximum for scientific applications. Besides scientific programming I have spent a significant amount of time on system programming for transputer machines (frequently because existing system software was inadequate for handling the scientific problems in which I was interested), and a lesser amount of time in system programming on SUN workstations running UNIX. I also have an interest in programming techniques and languages in general. It should also be said that I retain a strong interest in QCD.

References

[1] D.B. Carpenter, Two Dimensional Lattice QCD in Quenched Approximation. Nucl. Phys. B228 (1983) 365.

[2] D.B. Carpenter and C.F. Baillie, Free Fermion Propagators and Lattice Finite-Size Effects. Nucl. Phys. B260 (1985) 103.

[3] D.B. Carpenter, Blocking Transformations for Lattice Fermions. DESY Preprint 85-064. Unpublished.

[4] S.J. Hands and D.B. Carpenter, Lattice Sigma Model and Fermion Doubling. Nucl. Phys. B266 (1986) 285.

[5] C.R. Askew, D.B. Carpenter, J.T. Chalker, A.J.G. Hey, D.A. Nicole and D.J. Pritchard, Simulation of Statistical Mechanical Systems on Transputer Arrays. Computer Physics Communications 42 (1986) 21.

[6] C.R. Askew, D.B. Carpenter, J.T. Chalker, A.J.G. Hey, M. Moore, D.A. Nicole and D.J. Pritchard, Monte Carlo Simulation on Transputer Arrays. Parallel Computing 6 (1988) 247.

[7] J.R.A. Allwright and D.B. Carpenter, A Distributed Implementation of Simulated Annealing for the Travelling Salesman Problem. Parallel Computing 10 (1989) 335.

[8] D.B. Carpenter and J.T. Chalker, The Phase Diagram of a Generalised XY Model. J. Phys: Condens. Matter I (1989) 4907.

[9] B. Carpenter, C. Michael and M.J. Teper, O^{+} and 2^{+} Glueball Masses from Large Lattices in SU(2) Lattice Gauge Theory. Phys. Lett. B198 (1987) 511.

[10] C.R. Allton, D.B. Carpenter, V. Lubicz, L. Maiani, G. Martinelli and C.T. Sachrajda, Decay Constants of Heavy Mesons. Nuclear Physics B (Proc. Suppl.) 20 (1991) 504.

[11] A. Abada, C.R. Allton, Ph. Boucard, D.B. Carpenter, M. Crisafulli, J. Garland, S. Güsken, G. Martinelli, O. Pène, C.T. Sachrajda, R. Sarno, K. Schilling and R. Sommer Meson Spectroscopy and Decay Constants with Wilson Fermions at β = 6.4 Preprint (CERN-TH.6271/91, LPTHE-ORSAY 91/36, SHEP 91/92-3).